What do you understand by Hooke’s law of elasticity? Describe an experiment to determine Young’s modulus of an elastic material.
1
Expert's answer
2020-11-02T09:26:18-0500

Hooke's law is a statement according to which the deformation that occurs in an elastic body is proportional to the force applied to this body.

Take a spring, measure its length in the unstretched state, measure its diameter and find the cross-sectional area using the formula "S=\pi\times (\frac {d}{2})^2" , using a dynamometer, stretch the spring and write down the force and the new length of the spring and find the elongation, using the formula "E=\frac{\frac{F}{S}}{\frac {\Delta L}{L}}" find the young's modulus.